# Density-dependent sodium-storage mechanisms in hard carbon materials

### v1

@@ II.3 Hybrid Monte Carlo/molecular dynamics atomistic simulations
for sodium insertion and removal, respectively. $\Delta E$ is the energy difference between the trial and current configurations, $\beta={1}/({k_{\mathrm{B}}T})$, $V$ is the system’s volume, $N$ is the number of sodium atoms before the move, and $\lambda$ is the de Broglie thermal wavelength. A trial configuration is accepted if the corresponding acceptance probability is greater than a random number $r\in[0,1]$.
@@
We run MC/MD simulations for 100,000 steps for each structure for a chemical potential of 1.00-1.00, 1.24-1.24, 1.50-1.50 eV at 300 K. The chemical potential of 1.24-1.24 eV is related to the bulk sodium cohesive energy. Figure 3 shows the evolution of sodium concentration and dimensional change, defined as the relative change in the volume from the hard carbon volume after inserting sodium ions. All simulations reached the equilibrium after 50,000 MC steps, except for the low density structure at the highest chemical potential (i.e., 1.00-1.00 eV). The Monte Carlo simulations demonstrate a strong dependence of sodium storage on the structural density of nanoporous carbon. At low density (0.7 g cm-3), the framework accommodates the highest sodium concentration, reflecting the abundance of adsorption sites and the ease of pore filling. As the density increases to 1.6 g cm-3, the equilibrium concentration decreases significantly, suggesting that pore accessibility and connectivity become restricted. At the highest density investigated (1.9 g cm-3), sodium uptake is limited to 15%, which points to intercalation into a reduced number of energetically favorable sites rather than extended adsorption. This density-dependent trend is pronounced at $\mu=-1.00$ and $\mu=-1.24$ eV. In contrast, for $\mu=-1.50$ eV, sodium uptake is only weakly affected by density, indicating that in this regime insertion is governed primarily by the strong chemical driving force rather than by structural accessibility.
@@
The dimensional change of the carbon framework follows a similar trend. At low density ($\rho=0.7$ g cm-3) and high chemical potential, sodium insertion induces extreme swelling (up to 150 %), directly correlated with the large Na concentration. In this case, drastic structural rearrangements occur, including the breaking of sp3 bonds to accommodate volume expansion. At $\mu=-1.24$ eV, the dimensional change is much lower (13 %), suggesting a less disruptive insertion mechanism. In the intermediate density range (1.01.61.0–1.6 g cm-3) expansion remains moderate (9 and 6 % for 1.00-1.00 and 1.24-1.24 eV respectively), pointing to a mixed regime where adsorption still contributes but is increasingly restricted by pore confinement. At high density (1.9 g cm-3), dimensional changes strongly increase (102810–28 %), which can be attributed to the dominance of graphitic domains that undergo interlayer separation upon intercalation [Wang2021]. This resonates with the familiar result that Na ions do not intercalate in graphite and graphite-like materials that are able to accommodate the smaller Li ions. Consistent with the Na uptake results, the dimensional change at $\mu=-1.50$ eV remains largely independent of density for structures containing pores, underscoring the overriding influence of the chemical potential in this regime, in good agreement with an experimental dimensional change around 2 % [Alptekin2020, Escher2022] for a similar sodium concentration. The final sodiated structures of each simulation are available from Zenodo [Front2025_structures].
@@ III.1 Electrochemical properties
At $\mu=-1.00$ eV, the structures exhibit high storage capacities exceeding 3000 mAh g-1 in the extreme limit (which is unphysical). However, most of this uptake occurs at voltage approaching zero, which is unfavorable for practical operation due to the risk of sodium plating [Bommier2015]. As chemical potential decreases, the accessible capacity is progressively reduced, reaching $\approx 1000$ mAh g-1 at $\mu=-1.24$ eV and $\approx 200$ mAh g-1 at $\mu=-1.50$ eV. This trend illustrates the delicate balance between the driving force for Na insertion and the stability of the host framework. The role of density is also apparent across all values of $\mu$. Higher-density structures consistently yield slightly higher voltages at low degrees of sodiation, indicating stronger stabilization of sodium within compact host environments. In contrast, lower-density systems can accommodate significantly more sodium but at lower insertion voltages because, the larger the filled pore, the more the Na within resembles bulk metallic Na. All voltage curves exhibit a sharp initial drop within the first $50$ mAh g-1 or so, associated with the filling of the most favorable adsorption sites, followed by a gradual decay as insertion proceeds into weaker binding environments.
@@
To rationalize these trends, Fig. 5 decomposes the cumulative storage capacity into three distinct contributions: adsorption, intercalation, and pore filling, and illustrates the three corresponding classes of storage sites. Adsorption sites refer to Na-binding sites located on the carbon surface. These surface binding sites primarily consist of six-membered rings and edge motifs, with additional contributions from five- and seven-membered rings (Fig. 5(c)). Intercalation sites, in contrast, correspond to Na atoms residing in interlayer regions sandwiched between two graphene sheets (Fig. 5(d)). This type of binding site closely resembles those observed in the Li–graphite intercalation mechanism [Stevens2001]. Filling sites correspond to sodium atoms aggregated within nanopores or voids in the carbon matrix, where Na-Na interactions dominate over Na-C bonding (Fig. 5(e)). In low-density structures (0.71.00.7-1.0 g cm-3) sodium uptake is initially dominated by adsorption, but at higher capacities pore filling becomes the prevailing mechanism, accounting for more than 80 % of the total storage. This explains the very large capacities, but also why the associated voltages collapse toward zero, as pore filling corresponds to weakly bound states. At intermediate densities (1.31.61.3–1.6 g cm-3), adsorption remains dominant at the onset, but intercalation contributes substantially across the entire range, while pore filling is progressively suppressed. Finally, in dense HC (1.9 g cm-3), pore filling is negligible and sodium storage arises almost exclusively from adsorption and intercalation which lead to a limited capacity (200 mAh g-1). The same analysis is reported in the supplementary materials for a chemical potential of 1.00-1.00 eV (Figure S2-(a) and S3) and 1.50-1.50 eV (Figure S2-(c) and S4).
@@ IV Summary and conclusions
We have developed a comprehensive methodology to investigate sodium insertion mechanisms in hard carbon. Using machine-learning potentials coupled to grand-canonical Monte Carlo simulations, we performed realistic atomistic simulations from structure generation to the evaluation of electrochemical properties across a density range from 0.7 to 1.9 g cm-3. This approach provides direct insight into the relationship between HC density, sodium storage mechanism, and electrochemical performance. Our results reveal that sodium uptake in HC proceeds through three distinct contributions: adsorption, intercalation, and pore filling, whose relative importance strongly depends on mass density. Low-density HC (0.71.00.7-1.0 g cm-3) achieves high capacities dominated by pore filling but operates near the potential of metallic sodium, introducing a risk of Na plating. High-density HC (1.9 g cm-3) confines Na storage to adsorption and intercalation sites, resulting in safer operating voltages but limited capacity. Intermediate-density HC (1.31.61.3-1.6 g cm-3) achieves the best trade-off, combining balanced capacity, moderate voltage, and minimal structural deformation. These findings provide a microscopic explanation for the experimentally observed sloping and plateau regions in voltage profiles and clarify the density-driven origin of the capacity–voltage trade-off in HC anodes. The insights gained here establish density engineering as a key design strategy for next generation sodium-ion batteries, enabling the rational optimization of pore architecture, mechanical stability, and electrochemical performance.

### v2

@@ II.3 Hybrid Monte Carlo/molecular dynamics atomistic simulations
for sodium insertion and removal, respectively. $\Delta E$ is the energy difference between the trial and current configurations, $\beta={1}/({k_{\mathrm{B}}T})$, $V$ is the system’s volume, $N$ is the number of sodium atoms before the move, and $\lambda$ is the de Broglie thermal wavelength. A trial configuration is accepted if the corresponding acceptance probability is greater than a random number $r\in[0,1]$. Sodium insertion was simulated across the full volume of each HC structure, without explicit consideration of percolation pathways. Consequently, some sodium ions may occupy isolated voids that would be inaccessible in a real electrochemical environment. Therefore, absolute capacities obtained from these simulations may be overestimated. Nonetheless, the results allow for comparative analysis of trends in sodium insertion behavior across different structural densities.
@@
We run MC/MD simulations for 100,000 steps for each structure for a chemical potential of 1.00-1.00, 1.24-1.24, 1.50-1.50 eV at 300 K. The chemical potential of 1.24-1.24 eV corresponds to the bcc bulk sodium cohesive energy. Figure 3 shows the evolution of sodium concentration and dimensional change, defined as the relative change in the volume from the hard carbon volume after inserting sodium ions. All simulations reached the equilibrium after 50,000 MC steps, except for the low density structure at the highest chemical potential (i.e., 1.00-1.00 eV). The Monte Carlo simulations demonstrate a strong dependence of sodium storage on the structural density of nanoporous carbon. At low density (0.7 g cm-3), the framework accommodates the highest sodium concentration, reflecting the abundance of adsorption sites and the ease of pore filling. As the density increases to 1.6 g cm-3, the equilibrium concentration decreases significantly, suggesting that pore accessibility and connectivity become restricted. At the highest density investigated (1.9 g cm-3), sodium uptake is limited to 15%, which points to intercalation into a reduced number of energetically favorable sites rather than extended adsorption. This density-dependent trend is pronounced at $\mu=-1.00$ and $\mu=-1.24$ eV. In contrast, for $\mu=-1.50$ eV, sodium uptake is only weakly affected by density, indicating that in this regime insertion is governed primarily by the strong chemical driving force rather than by structural accessibility.
@@
The dimensional change of the carbon framework follows a similar trend. At low density ($\rho=0.7$ g cm-3) and high chemical potential, sodium insertion induces extreme swelling (up to 150 %), directly correlated with the large Na concentration. In this case, drastic structural rearrangements occur, including the breaking of sp3 bonds to accommodate volume expansion. At $\mu=-1.24$ eV, the dimensional change is much lower (13 %), suggesting a less disruptive insertion mechanism. In the intermediate density range (1.01.61.0–1.6 g cm-3) expansion remains moderate (9 and 6 % for 1.00-1.00 and 1.24-1.24 eV respectively), pointing to a mixed regime where adsorption still contributes but is increasingly restricted by pore confinement. At high density (1.9 g cm-3), dimensional changes strongly increase (102810–28 %), which can be attributed to the dominance of graphitic domains that undergo interlayer separation upon intercalation [47]. This resonates with the familiar result that Na ions do not intercalate in graphite and graphite-like materials that are able to accommodate the smaller Li ions. Consistent with the Na uptake results, the dimensional change at $\mu=-1.50$ eV remains largely independent of density for structures containing pores, underscoring the overriding influence of the chemical potential in this regime, in good agreement with an experimental dimensional change around 2 % [1, 14] for a similar sodium concentration. The final sodiated structures of each simulation are available from Zenodo [16].
@@ III.1 Electrochemical properties
At $\mu=-1.00$ eV, the structures exhibit high storage capacities approaching 4000 mAh g-1 in the extreme limit (which is unphysical). However, most of this uptake occurs at voltage approaching zero, which is unfavorable for practical operation due to the risk of sodium plating [4]. As chemical potential decreases, the accessible capacity is progressively reduced, reaching $\approx 1400$ mAh g-1 at $\mu=-1.24$ eV and $\approx 250$ mAh g-1 at $\mu=-1.50$ eV. This trend illustrates the delicate balance between the driving force for Na insertion and the stability of the host framework. The role of density is also apparent across all values of $\mu$. Higher-density structures consistently yield slightly higher voltages at low degrees of sodiation, indicating stronger stabilization of sodium within compact host environments. In contrast, lower-density systems can accommodate significantly more sodium but at lower insertion voltages because, the larger the filled pore, the more the Na within resembles bulk metallic Na. All voltage curves exhibit a sharp initial drop within the first $50$ mAh g-1 or so, associated with the filling of the most favorable adsorption sites, followed by a gradual decay as insertion proceeds into weaker binding environments.
@@
To rationalize these trends, Fig. 5 decomposes the cumulative storage capacity into three distinct contributions: adsorption, intercalation, and pore filling, and illustrates the three corresponding classes of storage sites. To quantify each contributions, we analysed the local atomic environment. Adsorption sites refer to Na-binding sites located on the carbon surface. These surface binding sites primarily consist of six-membered rings and edge motifs, with additional contributions from five- and seven-membered rings (Fig. 5(c)). Intercalation sites, in contrast, correspond to Na atoms residing in interlayer regions sandwiched between two graphene sheets (Fig. 5(d)). This type of binding site closely resembles those observed in the Li–graphite intercalation mechanism [40]. Filling sites correspond to sodium atoms aggregated within nanopores or voids in the carbon matrix, where Na-Na interactions dominate over Na-C bonding (Fig. 5(e)). In low-density structures (0.71.00.7-1.0 g cm-3) sodium uptake is initially dominated by adsorption, but at higher capacities pore filling becomes the prevailing mechanism, accounting for more than 80 % of the total storage. This explains the very large capacities, but also why the associated voltages collapse toward zero, as pore filling corresponds to weakly bound states. At intermediate densities (1.31.61.3–1.6 g cm-3), adsorption remains dominant at the onset, but intercalation contributes substantially across the entire range, while pore filling is progressively suppressed. Finally, in dense HC (1.9 g cm-3), pore filling is negligible and sodium storage arises almost exclusively from adsorption and intercalation which lead to a limited capacity (250 mAh g-1). The same analysis is reported in the supplementary materials for a chemical potential of 1.00-1.00 eV (Figure S3-(a) and S4) and 1.50-1.50 eV (Figure S3-(c) and S5).
@@ IV Summary and conclusions
We have developed a comprehensive methodology to investigate sodium insertion mechanisms in hard carbon. Using machine-learning potentials coupled to grand-canonical Monte Carlo simulations, we performed realistic atomistic simulations from structure generation to the evaluation of electrochemical properties across a density range from 0.7 to 1.9 g cm-3. This approach provides direct insight into the relationship between HC porosity, sodium storage mechanism, and electrochemical performance. Our results reveal that sodium uptake in HC proceeds through three distinct contributions: adsorption, intercalation, and pore filling, whose relative importance strongly depends on mass density. Low-density HC (0.71.00.7-1.0 g cm-3) achieves high capacities dominated by pore filling but operates near the potential of metallic sodium, introducing a risk of Na plating. High-density HC (1.9 g cm-3) confines Na storage to adsorption and intercalation sites, resulting in safer operating voltages but limited capacity. Intermediate-density HC (1.31.61.3-1.6 g cm-3) achieves the best trade-off, combining balanced capacity, moderate voltage, and minimal structural deformation. These findings provide a microscopic explanation for the experimentally observed sloping and plateau regions in voltage profiles and clarify the density-driven origin of the capacity–voltage trade-off in HC anodes. The insights gained here establish density engineering as a key design strategy for next generation sodium-ion batteries, enabling the rational optimization of pore architecture, mechanical stability, and electrochemical performance.